OpenStudy (anonymous):

What is the 10th term in the geometric sequence: 32, 16, 8, 4...?

OpenStudy (anonymous):

a10 = a1*(r^(n-1) ) here a1=32 and r=16/32=8/16=4/8 ......=1/2 and n=10 and so on....